For a chi-square test of independence for this contingency table, what is the number of degrees of freedom?

You observe 100 randomly selected college students to find out whether they arrive on time or late for their classes. The table below gives a two-way classification for these students.

GenderOn TimeLate
Female359
Male4313


For a chi-square test of independence for this contingency table, what is the number of degrees of freedom?
Answer #1

Degrees of freedom = (no of rows - 1) * (no of cols - 1) = (2 - 1) * (2 - 1) = 1 (ans)
Answer #2

Solution from giver dabu Ayung Cured Nol Curred. 10 LI - 59 Degrees of freedom = (rown) * (corumn-1) = (2-1) ok (2-1) Degrees